# On the knot types of periodic Reeb orbits of dynamically convex contact forms

### v1

@@ 1. Introduction and main results
Knotted structures in the dynamics of flows in dimension three have been studied since the XIXth century with the works of Lord Kelvin and Helmholtz. The task of understanding knots of periodic orbits is a topic that was visited by several authors during the XXth century. For example, in a foundational paper [BiWill] Birman and Williams prove that periodic orbits in the Lorentz attractor are all fibered knots. Here we consider Reeb flows on closed contact $3$-manifolds and study the transverse knot types realized by their closed orbits.
As observed by Etnyre and Ghrist [EG], one may take the complexity of the transverse knots realized by the closed orbits as a rough measure of the topological complexity of the flow. In the following discussion consider, for simplicity, the case of the standard contact $3$-sphere $(S^{3},\xi_{0})$, where $S^{3}=\{(z_{0},z_{1})\in\mathbb{C}^{2}:|z_{0}|^{2}+|z_{1}|^{2}=1\}$ and $\xi_{0}=TS^{3}\cap iTS^{3}$. One may ask how complex a Reeb flow on $(S^{3},\xi_{0})$ can be from this viewpoint. The result from [EG] asserts that there exist real-analytic examples where all transverse knot types are simultaneously realized. The main goal of this paper is to study concrete geometric properties that prevent this phenomenon from happening, and force restrictions on the transverse knot types that can arise as closed orbits. Specifically, we look at the property of dynamical convexity, an important concept introduced by Hofer, Wysocki and Zehnder (HWZ) in [convex]. The precise definition will be recalled below. Two rich sources of examples are the strictly convex energy levels in a symplectic vector space, see Theorem 3.4 in  [convex], and the unit cotangent bundles of Finsler two-spheres with reversibility $r\geq 1$ and flag curvatures pinched by strictly more than $(r/(r+1))^{2}$, see [HP, HS_Cambridge].
@@
See § 2.4 for a precise definition of the Conley-Zehnder index relative to a capping disk.
@@
Let $(W,\omega)$ be a connected compact symplectic manifold with contact-type boundary $(V,\lambda)$. Suppose that $(W,\omega)$ is symplectically aspherical, i.e. $\int_{S^{2}}f^{*}\omega=0$ for every smooth map $f:S^{2}\to W$. Assume that some connected component $V_{0}^{+}\subset V^{+}$ has a simple periodic orbit $P$ with a symplectic slicing disk $D$ such that $(V_{0}^{+},\lambda)$ is dynamically convex in $W$ relative to $(P,D)$. Then either there is a contractible periodic Reeb orbit on $V^{-}$ with period less than $\int_{D}\omega$, or all of the following hold:
@@
A source of symplectically null-homologous transverse links are the transverse intersections of a complex curve in $\mathbb{C}^{2}$ with the round $3$-sphere. Using results of Micallef and White [MW] we can perturb the complex structure to an almost complex structure, and the curve to a pseudo-holomorphic curve with only transverse double points. These can be swapped for genus using a standard trick, and we thus get an embedded symplectic surface spanning the link. Such links are examples of what are known as transverse $\mathbb{C}$-links. The general definition allows for transverse intersections of a complex curve with a strictly pseudo-convex $3$-sphere bounding a Stein $4$-ball in $\mathbb{C}^{2}$, see Definition 4.74 in [R2]. But according to Proposition 4.75 in [R2], it was observed by Boileau and Orevkov in [BO] that results of Eliashberg [E] imply that all transverse $\mathbb{C}$-links can be realised as intersections of complex curves with the round $3$-sphere. Moreover, Rudolph proved in [R1] that every so-called quasipositive link is smoothly isotopic to a transverse $\mathbb{C}$-link, and therefore smoothly isotopic to a symplectically null-homologous link. The converse is proved in [BO], namely, every transverse $\mathbb{C}$-link is quasipositive or, more generally, every symplectically null-homologous link on the boundary of a symplectic $4$-ball is quasipositive.
@@
Acknowledgments. U.H. acknowledges the support by the RWTH Aachen, the DFG SFB/TRR 191 ‘Symplectic Structures in Geometry, Algebra and Dynamics’, Projektnummer 281071066-TRR 191, and FAPERJ Grant ‘Pesquisador Visitante’ 203.664/2025. PS was partially supported by the National Natural Science Foundation of China (grant number W2431007). PS thanks the support of the Shenzhen International Center for Mathematics - SUSTech.
@@ 2.7. Fredholm indices and moduli spaces
An important consequence is that we have automatic transversality, i.e. $D_{\widetilde{u}}$ at an embedded fast plane $\widetilde{u}$ is always a surjective Fredholm operator. This can be proved by an indirect argument as follows. There is no loss of generality to deform the normal bundle so that it coincides with $u^{*}\xi$ over $\mathbb{C}\setminus B_{R}(0)$, $R\gg 1$. A global trivializing section of the normal bundle then induces, up to homotopy, a $d\lambda$-symplectic trivialization $\sigma_{N}$ of $x(T\cdot)^{*}\xi$ which winds +1+1 with respect to $\sigma$. This means that $\operatorname{wind}_{\sigma_{N}}(\mu)=\operatorname{wind}_{\sigma}(\mu)-1$ for every eigenvalue $\mu$ of the asymptotic operator. Moreover, a nontrivial section $\zeta\in\ker D_{\widetilde{u}}$ admits an asymptotic behavior governed by an eigensection of the asymptotic operator associated to an eigenvalue $\nu<\delta$, see [fast, Theorem 6.1] or [sie_CPAM, Theorem A.1]. Hence, $\zeta$ does not vanish near $\infty$ and the total algebraic count of zeros of $\zeta$ is equal to $\operatorname{wind}_{\sigma_{N}}(\nu)=\operatorname{wind}_{\sigma}(\nu)-1\leq\alpha^{<\delta}_{\sigma}(P)-1=1-1=0$. But the equation $D_{\widetilde{u}}\zeta=0$ allows us to use Carleman’s similarity principle to say that zeros are isolated and count positively. The important conclusion that $\zeta$ never vanishes. Since the Fredholm index is $2$, we would find $3$ linearly independent sections of the kernel if the linearization is not surjective. But the normal bundle is two-dimensional, hence a nontrivial linear combination of them would have to vanish at some point, contradiction.
@@
It turns out that, under our standing assumption that $\lambda$ is nondegenerate, the topologies on $\mathcal{M}^{\rm fast}(P,\bar{J})$ and on $\mathcal{M}^{\rm fast}_{1}(P,\bar{J})$ inherited from the functional analytic set-up used for the Fredholm theory coincides with the topology of $C^{\infty}_{\rm loc}$-convergence. There are situations where this can be proved dropping the assumption that $\lambda$ is nondegenerate [hryn_jsg, elliptic].
@@ 2.8. Intersection theory
Write $\widetilde{u}_{n}(s,t)$ instead of $\widetilde{u}_{n}(e^{2\pi(s+it)})$, and similarly $\widetilde{u}(s,t)$ for $s$ large enough. Let $\mathcal{N}$ be a small tubular neighborhood of $x(\mathbb{R})$ in $V^{+}$. We can assume, with no loss of generality, that $\sigma$ extends to a $d\lambda$-symplectic trivializing frame of $\xi|_{\mathcal{N}}$ which we still denote by $\sigma$. It follows from a suitable application of the Monotonicity Lemma [sft_comp, Lemma 5.2] together with results on cylinders with small contact area from [small_area] that there exists $s_{0}\gg 1$ such that
@@ 2.10. SFT-limits of fast planes
Let $\widetilde{u}_{0},S_{0},j_{0},\Gamma^{+}_{0},\Gamma^{-}_{0},D_{0}]$ be the nodal curve at the level zero of $\mathbf{u}$. If $D_{0}\neq\emptyset$, $k_{+}=k_{-}=0$ then there exists nonconstant $\bar{J}$-sphere.
@@
If $k_{-}\geq 1$ then there exists a periodic Reeb orbit $P_{*}=(x_{*},T_{*})$ in $(V^{-},\lambda)$ which is contractible in $V^{-}$ and satisfies
@@
In any case, choose $j_{*}\in\{1,\dots,h\}$ with the following property: $c_{j_{*}}$ corresponds to a negative puncture of a component of $\widetilde{u}_{0}$ that either has a positive puncture, or has no positive punctures but is connected to a component of $\widetilde{u}_{0}$ with a positive puncture by a nodal pair. Note that in the latter case there is exactly one such nodal due to the fact that $S^{{\bf u},r}$ has genus zero. Such $j_{*}$ must exist, since otherwise $\widetilde{u}_{0}$ has no positive puncture, contradicting the fact that the entire building is the SFT-limit of a sequence of planes in $\overline{W}$ with one positive puncture. Let $P_{*}=(x_{*},T_{*})$ be the asymptotic limit of $\widetilde{u}_{0}$ at the negative puncture corresponding to $c_{j_{*}}$.
@@ 2.11. The symplectic capping disk as a finite-energy plane
To prove surjectivity of (32) we will show that ${\rm ev}(\mathcal{M}^{\rm fast}_{1}(P,\bar{J},e))$ is closed in $\overline{W}$. To this end suppose that $\widetilde{v}_{n},z_{n}]\in\mathcal{M}^{\rm fast}_{1}(P,\bar{J},e)$ and that ${\rm ev}[\widetilde{v}_{n},z_{n}]$ converges to a point $q\in\overline{W}$. Up to choice of a subsequence, we may assume that $\widetilde{v}_{n},\mathbb{C}\cup\{\infty\},i,\{\infty\},\emptyset,\emptyset]$ SFT-converges to a building $\mathbf{u}$ of height $k_{-}|1|k_{+}$. We claim that $k_{+}=0$. If not then Lemma 2.15 implies that $V_{0}^{+}$ is not dynamically convex in $W$ relative to $(P,D)$. If $k_{-}\geq 1$ then Lemma 2.17 provides a contractible periodic Reeb orbit $P_{*}=(x_{*},T_{*})$ in $V^{-}$ such that $T_{*}$ satisfies (19) for all $b>0$. The monotonicity lemma implies that the term $\int_{\widetilde{u}_{0}^{-1}(W)}\widetilde{u}_{0}^{*}\overline{\omega}_{b}$ has a positive lower bound independent of $b$, where $\widetilde{u}_{0}$ is the map representing the level zero of $\mathbf{u}$. Taking the limit as $b\to 0^{+}$ we get $T_{*}<T$. This is one of the alternatives in Theorem 1.4, and we are now left with the case $k_{-}=0$. In this case, $\mathbf{u}$ consists of a single nodal curve $\widetilde{u}_{0}$. Lemma 2.16 tells us that there is a $\bar{J}$-holomorphic sphere if $\widetilde{u}_{0}$ has nodal points, but such spheres are excluded since $(W,\omega)$ is aspherical. Hence $\widetilde{u}_{0}$ is a finite-energy plane asymptotic to $P$. It is the $C^{\infty}_{\rm loc}$-limit of a suitable reparametrization of the planes in the sequence $\widetilde{v}_{n}$ which we still denote by $\widetilde{v}_{n}$. Lemma 2.11 implies that $\widetilde{u}_{0}$ represents a curve in $\mathcal{M}^{\rm fast}(P,\bar{J},e)$. SFT convergence implies that $\widetilde{v}_{n},z_{n}]$ converges in $\mathcal{M}^{\rm fast}_{1}(P,\bar{J},e)$ to $\widetilde{u}_{0},\zeta]$ for some $\zeta\in\mathbb{C}$. Hence ${\rm ev}[\widetilde{u}_{0},\zeta]=q$.

### v2

@@ 1. Introduction and main results
Knotted structures in the dynamics of flows in dimension three have been studied since the XIXth century with the works of Lord Kelvin and Helmholtz. The task of understanding knots of periodic orbits is a topic that was visited by several authors during the XXth century. For example, in a foundational paper [BiWill] Birman and Williams prove that all periodic orbits in the Lorentz attractor are fibered knots. More recently, it was proved in [AF] that the same is true for every periodic orbit of the geodesic flow of a Riemannian two-sphere with an explicit pinching condition on the curvatures. Here we consider Reeb flows on closed contact $3$-manifolds and study the transverse knot types realized by their closed orbits.
As observed by Etnyre and Ghrist [EG], one may take the complexity of the transverse knots realized by the closed orbits as a rough measure of the topological complexity of the flow. In the following discussion consider, for simplicity, the case of the standard contact $3$-sphere $(S^{3},\xi_{0})$, where $S^{3}=\{(z_{0},z_{1})\in\mathbb{C}^{2}:|z_{0}|^{2}+|z_{1}|^{2}=1\}$ and $\xi_{0}=TS^{3}\cap iTS^{3}$. One may ask how complex a Reeb flow on $(S^{3},\xi_{0})$ can be from this viewpoint. The result from [EG] asserts that there exist real-analytic examples where all transverse knot types are simultaneously realized. The main goal of this paper is to study concrete geometric properties that prevent this phenomenon from happening, and force restrictions on the transverse knot types that can arise as closed orbits. Specifically, we look at the property of dynamical convexity, an important concept introduced by Hofer, Wysocki and Zehnder (HWZ) in [convex]. The precise definition will be recalled below. Two rich sources of examples are the strictly convex energy levels in a symplectic vector space, see Theorem 3.4 in [convex], and the unit cotangent bundles of Finsler two-spheres with reversibility $r\geq 1$ and flag curvatures pinched by strictly more than $(r/(r+1))^{2}$, see [HP, HS_Cambridge]. Additionally, dynamical convexity has been observed in numerous celestial mechanics models [dPKSS2024, JK, LS2025, HLOSY2025].
@@
See § 2.4 for a precise definition of the Conley-Zehnder index relative to a capping disk. Motivated by results of Geiges and Zehmisch [GZ], we encode our analysis in the form of a symplectic characterization result for star-shaped domains in $(\mathbb{C}^{2},\omega_{0})$.
@@
Let $(W,\omega)$ be a connected compact symplectic manifold with contact-type boundary $(V,\lambda)$. Suppose that $(W,\omega)$ is symplectically aspherical, i.e. $\int_{S^{2}}f^{*}\omega=0$ for every smooth map $f:S^{2}\to W$. Assume that some connected component $V_{0}^{+}\subset V^{+}$ has a simple periodic orbit $P$ with a symplectic slicing disk $D$ such that $(V_{0}^{+},\lambda)$ is dynamically convex in $W$ relative to $(P,D)$. Then either there is a periodic Reeb orbit on $V^{-}$ with period less than $\int_{D}\omega$, or all of the following hold:
@@
If $(W,\omega)$ is an exact symplectic cobordism, then the first alternative in the conclusions of Theorem 1.4 can be improved to provide a contractible periodic Reeb orbit in $V_{-}$.
Remark 1.7.
@@
A source of symplectically null-homologous transverse links are the transverse intersections of a complex curve in $\mathbb{C}^{2}$ with the round $3$-sphere. Using results of Micallef and White [MW] we can perturb the complex structure to an almost complex structure, and the curve to a pseudo-holomorphic curve with only transverse double points. These can be swapped for genus using a standard trick, and we thus get an embedded symplectic surface spanning the link. Such links are examples of what are known as transverse $\mathbb{C}$-links. The general definition allows for transverse intersections of a complex curve with a strictly pseudo-convex $3$-sphere bounding a Stein $4$-ball in $\mathbb{C}^{2}$, see Definition 4.74 in [R2]. But according to Proposition 4.75 in [R2], it was observed by Boileau and Orevkov in [BO] that results of Eliashberg [E] imply that all transverse $\mathbb{C}$-links can be realized as intersections of complex curves with the round $3$-sphere. Moreover, Rudolph proved in [R1] that every so-called quasipositive link is smoothly isotopic to a transverse $\mathbb{C}$-link, and therefore smoothly isotopic to a symplectically null-homologous link. The converse is proved in [BO], namely, every transverse $\mathbb{C}$-link is quasipositive or, more generally, every symplectically null-homologous link on the boundary of a symplectic $4$-ball is quasipositive.
@@
Acknowledgments. We would like to warmly thank Shah Faisal for spotting a mistake in the statement Theorem 1.4 of a previous version of this paper: the periodic orbit claimed to exist in $V_{-}$ is not necessarily contractible. U.H. acknowledges the support from the RWTH Aachen, the SFB/TRR 191 ‘Symplectic Structures in Geometry, Algebra and Dynamics’, funded by the DFG (Projektnummer 281071066–TRR 191), and the FAPERJ Grant ‘Pesquisador Visitante’ 203.664/2025. PS was partially supported by the National Natural Science Foundation of China (grant number W2431007). PS thanks the support of the Shenzhen International Center for Mathematics - SUSTech. RS gratefully acknowledges financial support from the DFG grant BR 5251/1-1, and from the SFB/TRR 191 “Symplectic Structures in Geometry, Algebra and Dynamics” funded by the DFG (Projektnummer 281071066–TRR 191) during the completion of this project.
@@ 2.7. Fredholm indices and moduli spaces
An important consequence is that we have automatic transversality, i.e. $D_{\widetilde{u}}$ at an embedded fast plane $\widetilde{u}$ is always a surjective Fredholm operator. This can be proved by an indirect argument as follows. There is no loss of generality to deform the normal bundle so that it coincides with $u^{*}\xi$ over $\mathbb{C}\setminus B_{R}(0)$, $R\gg 1$. A global trivializing section of the normal bundle then induces, up to homotopy, a $d\lambda$-symplectic trivialization $\sigma_{N}$ of $x(T\cdot)^{*}\xi$ which winds +1+1 with respect to $\sigma$. This means that $\operatorname{wind}_{\sigma_{N}}(\mu)=\operatorname{wind}_{\sigma}(\mu)-1$ for every eigenvalue $\mu$ of the asymptotic operator. Moreover, a nontrivial section $\zeta\in\ker D_{\widetilde{u}}$ admits an asymptotic behavior governed by an eigensection of the asymptotic operator associated to an eigenvalue $\nu<\delta$, see Theorem 6.1 in [fast] or Theorem A.1 in [sie_CPAM]. Hence, $\zeta$ does not vanish near $\infty$ and the total algebraic count of zeros of $\zeta$ is equal to $\operatorname{wind}_{\sigma_{N}}(\nu)=\operatorname{wind}_{\sigma}(\nu)-1\leq\alpha^{<\delta}_{\sigma}(P)-1=1-1=0$. But the equation $D_{\widetilde{u}}\zeta=0$ allows us to use Carleman’s similarity principle to say that zeros are isolated and count positively. The important conclusion that $\zeta$ never vanishes. Since the Fredholm index is $2$, we would find $3$ linearly independent sections of the kernel if the linearization is not surjective. But the normal bundle is two-dimensional, hence a nontrivial linear combination of them would have to vanish at some point, contradiction.
@@
It turns out that, under our standing assumption that $\lambda$ is nondegenerate, the topologies on $\mathcal{M}^{\rm fast}(P,\bar{J})$ and on $\mathcal{M}^{\rm fast}_{1}(P,\bar{J})$ inherited from the functional analytic set-up used for the Fredholm theory coincides with the topology of $C^{\infty}_{\rm loc}$-convergence. There are situations where this can be proved dropping the assumption that $\lambda$ is nondegenerate [hryn_jsg],[elliptic].
@@ 2.8. Intersection theory
Write $\widetilde{u}_{n}(s,t)$ instead of $\widetilde{u}_{n}(e^{2\pi(s+it)})$, and similarly $\widetilde{u}(s,t)$ for $s$ large enough. Let $\mathcal{N}$ be a small tubular neighborhood of $x(\mathbb{R})$ in $V^{+}$. We can assume, with no loss of generality, that $\sigma$ extends to a $d\lambda$-symplectic trivializing frame of $\xi|_{\mathcal{N}}$ which we still denote by $\sigma$. It follows from a suitable application of the Monotonicity Lemma, Lemma 5.2 in [sft_comp] together with results on cylinders with small contact area from [small_area] that there exists $s_{0}\gg 1$ such that
@@ 2.10. SFT-limits of fast planes
Let $\widetilde{u}_{0},S_{0},j_{0},\Gamma^{+}_{0},\Gamma^{-}_{0},D_{0}]$ be the nodal curve at the level zero of $\mathbf{u}$. If $D_{0}\neq\emptyset$, $k_{+}=k_{-}=0$ then there exists non-constant $\bar{J}$-sphere.
@@
If $k_{-}\geq 1$ then there exists a periodic Reeb orbit $P_{*}=(x_{*},T_{*})$ in $(V^{-},\lambda)$ satisfying
@@
In any case, choose $j_{*}\in\{1,\dots,h\}$ with the following property: $c_{j_{*}}$ corresponds to a negative puncture of a component of $\widetilde{u}_{0}$ that either has a positive puncture, or has no positive punctures but is connected to a component of $\widetilde{u}_{0}$ with a positive puncture by nodal pairs. Note that in the latter case there is exactly one such nodal pair due to the fact that $S^{{\bf u},r}$ has genus zero and that $(W,\omega)$ is symplectically aspherical. Such $j_{*}$ must exist, since otherwise $\widetilde{u}_{0}$ has no positive puncture, contradicting the fact that the entire building is the SFT-limit of a sequence of planes in $\overline{W}$ with one positive puncture. Let $P_{*}=(x_{*},T_{*})$ be the asymptotic limit of $\widetilde{u}_{0}$ at the negative puncture corresponding to $c_{j_{*}}$.
@@ 2.11. The symplectic capping disk as a finite-energy plane
To prove surjectivity of (32) we will show that ${\rm ev}(\mathcal{M}^{\rm fast}_{1}(P,\bar{J},e))$ is closed in $\overline{W}$. To this end suppose that $\widetilde{v}_{n},z_{n}]\in\mathcal{M}^{\rm fast}_{1}(P,\bar{J},e)$ and that ${\rm ev}[\widetilde{v}_{n},z_{n}]$ converges to a point $q\in\overline{W}$. Up to choice of a subsequence, we may assume that $\widetilde{v}_{n},\mathbb{C}\cup\{\infty\},i,\{\infty\},\emptyset,\emptyset]$ SFT-converges to a building $\mathbf{u}$ of height $k_{-}|1|k_{+}$. We claim that $k_{+}=0$. If not then Lemma 2.15 implies that $V_{0}^{+}$ is not dynamically convex in $W$ relative to $(P,D)$. If $k_{-}\geq 1$ then Lemma 2.17 provides a periodic Reeb orbit $P_{*}=(x_{*},T_{*})$ in $V^{-}$ such that $T_{*}$ satisfies (19) for all $b>0$. The monotonicity lemma implies that the term $\int_{\widetilde{u}_{0}^{-1}(W)}\widetilde{u}_{0}^{*}\overline{\omega}_{b}$ has a positive lower bound independent of $b$, where $\widetilde{u}_{0}$ is the map representing the level zero of $\mathbf{u}$. Taking the limit as $b\to 0^{+}$ we get $T_{*}<T$. This is one of the alternatives in Theorem 1.4, and we are now left with the case $k_{-}=0$. In this case, $\mathbf{u}$ consists of a single nodal curve $\widetilde{u}_{0}$. Lemma 2.16 tells us that there is a $\bar{J}$-holomorphic sphere if $\widetilde{u}_{0}$ has nodal points, but such spheres are excluded since $(W,\omega)$ is aspherical. Hence $\widetilde{u}_{0}$ is a finite-energy plane asymptotic to $P$. It is the $C^{\infty}_{\rm loc}$-limit of a suitable reparametrization of the planes in the sequence $\widetilde{v}_{n}$ which we still denote by $\widetilde{v}_{n}$. Lemma 2.11 implies that $\widetilde{u}_{0}$ represents a curve in $\mathcal{M}^{\rm fast}(P,\bar{J},e)$. SFT convergence implies that $\widetilde{v}_{n},z_{n}]$ converges in $\mathcal{M}^{\rm fast}_{1}(P,\bar{J},e)$ to $\widetilde{u}_{0},\zeta]$ for some $\zeta\in\mathbb{C}$. Hence ${\rm ev}[\widetilde{u}_{0},\zeta]=q$.